OpenStudy (lolipop_drum312):

Math check help please! Solve for r. s=2πrh A. 2πsh=r B. s/2πh=r <-- C. s/2π=r D. s/2=r Solve for c. a(b − c)=d A. c=d − ab/a B. c=ab − d/b C. c=ab + d/a <-- D. c=ab − d/a

OpenStudy (lolipop_drum312):

Solve for w. P = 2l + 2w A. w=2l−P/2 B. w=P−2l/l C. w=P−l D. w=P−2l/2

OpenStudy (lolipop_drum312):

i think ^ is B

OpenStudy (igreen):

Your first one is correct.

OpenStudy (lolipop_drum312):

ok :)

OpenStudy (lolipop_drum312):

is the second wrong

OpenStudy (igreen):

The 3rd one is incorrect..

OpenStudy (lolipop_drum312):

oh then how do i solve for that one becuz my second choice would be d

OpenStudy (lolipop_drum312):

idk

OpenStudy (igreen):

The 2nd one is also incorrect.

OpenStudy (igreen):

Yes.. D is correct. P = 2l + 2w Subtract 2l to both sides: P - 2l = 2w Divide 2 to both sides: (P - 2l)/2 = w

OpenStudy (lolipop_drum312):

AH! OK! so how about 3